You are processing two prescriptions for a seven (7) year old. One is for Tylenol 160mg/5mL and the other one is for Motrin 100mg/5mL. The pediatric dosing of acetaminophen is 10 mg/kg/dose to 15 mg/kg/dose and for ibuprofen is 10 mg/kg/dose. The provider want the lowest dose for both medications. How many milliliters per dose of Tylenol and Motrin should be dispensed? The patient weighs 47 lbs. Show your calculations.
Answer #1

Firstly the weight of patient is 47 lb =21.36 kg

First. Tylenol contains paracetamol salt

the pediatric dose is =10 mg/kg/dose to  15 mg/kg/dose

as per weight required dose= 10 X 21.36= 213.6 mg to 15 X 21.36=320.4

mEnad dose range is 213.6 mg to 320.4 mg

Given that available Tylenol 160mg/5mL= 32 mg/ml

for 213.6 mg required milliliters of Tylenol as per dose= 213.6/32= 6.7 ml

for 320.4 mg equired milliliters of Tylenol as per dose=320.4/32=10.01 ml

means we can give 6.7 to 10.1 ml of of Tylenol as per dose

Second: Motrin contains ibuprofen as salt

the pediatric dose is =10 mg/kg/dose

as per weight required dose= 10 X 21.36= 213.6 mg

Given that available Motrin 100mg/5mL= 20 mg/ml

for 213.6 mg required milliliters of Motrin as per dose will be = 213.6/20=10. 7 ml